A jogger consistently runs at a pace where she can complete three miles in 27 minutes and 30 seconds. Assume the jogger can keep this pace indefinitely. Using Mathcad, develop a function to describe the distance the jogger travels given a specific length of time. (a) Determine the distance covered after 10 minutes of jogging, after 25 minutes of jogging, and after 40 minutes of jogging using your function. (b) Then, define a range variable to represent time from 0 to 120 minutes in increments of 5 minutes. Plot the discretized function against the range variable on two separate plots. On the first plot, express distance in the magnitudes associated with km and time in the magnitude associated with hr. On the second plot, express distance in the magnitudes associated with miles and time in the magnitudes associated with minutes.
Answer #1

Given,

Time = 27 minutes and 30 seconds = 27.5 minutes

Distance = 3 miles

Calculate speed = distannce/time

=>speed = 3/27.5 miles/min

or

=>(3*1.6)*(60/27.5) km/hr (1 miles=1.6 km approx. and 1 hr =60 mins)

(a).

Distance Covered = speed * specified time

=> Distance = (3/27.5) * t here, t is time in minutes.

(b).

Time range graph

For the first plot in km/hr, use distance(d) = (3*16)*(60/27.5)*t where t is time in hours

For the Second plot in km/hr, use distance(d) = (3/27.5)*t where t is time in minutes

Relations between distance and time can be depicted in graphs by the above functions.

The concepts I provided in the answer may be used to work in Mathcad